Find the slope and the y-intercept of the line. Write the fractional answer in lowest terms. y=-1/4x-8/5 The slope is ?
Found 2 solutions by solver91311, rfer:
Answer by solver91311(24713)



When you have a function in the form , the slope is the coefficient on and the -intercept is the constant term.


John


Answer by rfer(16322)

y=mx+b
m= slope=-1/4
b= y int=-8/5=-1 3/5